OpenStudy (solomonzelman):

Attempt 2: Can someone teach me how to master cubic functions? (Don't come just to troll)

OpenStudy (solomonzelman):

I know some cubic functions, but not the harder ones.

OpenStudy (anonymous):

give me an example of what you struggle with and ill try to help

OpenStudy (solomonzelman):

I am not, because, I'm only a senior. but anyway, I was stock with equations that look like \[f(x)=x^3+x+31\]

OpenStudy (anonymous):

well remember that when the degree is even the end arrows go in the same direction and if the degree is odd the endpoints will go in different directions. do you know how to find the zeros? or the remainder? do you remember synthetic division

OpenStudy (solomonzelman):

I know how to divide polynomials, find zeros, and etc, but this one is more complicated then (lets say) \[f(x)=x^3-4x^2-9x+36~~~~or~~something~~w/~~~2nd~~degree.\]

OpenStudy (anonymous):

same process. the degree is cubic so they go in diff directions and the coefficient is positive so the endpoints will be in Q1 and Q3 if it had 2nd degree its just a quadratic equation and that is just a parabola.

OpenStudy (solomonzelman):

I know this from trig, that degree is positive, then looks like..... (I learned this doing parabolas) but can you walk me through? How do I do \[f(x)=x^3+x+31~~~~~~?\]

OpenStudy (anonymous):

what are you wanting to find? the zeros? we know that there are 3 possible solutions to this the factors of the constant will be possible zeros. that is a really ugly factor on that one since the factors are decimels

OpenStudy (solomonzelman):

Well, lets find the zeros.

OpenStudy (anonymous):

hmm. other than using a graphing calculator i'm not sure with a prime constant that doesn't factor into the cubic function. sorry i couldn't help. if you factor it on wolfram it gives some ugly decimals http://www.wolframalpha.com/input/?i=zeros+of+x%5E3%2Bx%2B31

OpenStudy (solomonzelman):

Tnx for helping, but I want a formula-way to do it.

OpenStudy (anonymous):

im curious as well. I tried to find the derivative of it and solve the quadratic but that will only give maximae coordinates not zeros

OpenStudy (solomonzelman):

Yeah, also finding the derivatives wouldn't help, because I don't know what derivative is.

OpenStudy (anonymous):

well, just using the power rule if we had \[2x^3+2x^2+2x+2\] the derivative would be \[6x^2+4x+2\] and if you solve for the roots you would get the maxima on some problems notice the relations between the degrees and the coefficients

OpenStudy (anonymous):

sorry I couldn't help more. gotta get to class.

OpenStudy (anonymous):

alright. I asked my teacher and you have to plug numbers in that are consecutive integers until you get a sign change. once you find the two integers that create a sign change you start plugging in numbers between them to find where it =0. does that make sense?